# A Step to Decouple Optimization in 3DGS

### v1

@@ 1 Introduction
Novel view synthesis is a fundamental task in both computer graphics and computer vision. Owing to its highly parallelized design and efficient use of GPU, 3DGS (kerbl20233d, ) leverages explicit primitive-based representations to achieve significantly higher efficiency than implicit representation NeRF (mildenhall2020nerf, ), while still delivering competitive reconstruction quality. 3DGS sparks a wave of research and rapid extensions to various downstream applications (matsuki2024gaussian, ; chen2024text, ; dai2024high, ).
@@
As a representation directly optimized via backpropagation, 3DGS inherits the Adam (kinga2015method, ) from DNNs, and in default attributes of all primitives are updated simultaneously within each iteration, which we call synchronous optimization in this work. Given the viewpoint relationship, an ideal scenario is that the optimization of primitives is guided by gradients from visible viewpoints in different directions. However, upon closer examination, we observe that even when primitives are invisible under a given viewpoint and their gradients are set to zero, the update-step coupling induced by Adam and synchronous optimization implicitly rescales the optimizer state and causes updating in these invisible viewpoints. From the perspective of optimization efficiency, (mallick2024taming, ) introduces Sparse Adam, which restricts updates to primitives visible under viewpoints while filtering out those invisible with zero gradients, thereby enabling asynchronous optimization. Nevertheless, this optimizer has not been widely adopted, since its efficiency gains come at the expense of degraded performance. However, we think that such a viewpoint-related optimizer remains promising, with the current limitation lying in the insufficient recognition of update-step coupling.
@@
Furthermore, previous study (loshchilov2017decoupled, ) in DNNs points out the gradient coupling in Adam makes the regularization unstable. In current 3DGS, a variety of regularization losses (yu2024gaussian, ) are introduced to improve reconstruction quality or mitigate redundancy111Redundancy is strictly defined as minimizing active primitives without sacrificing reconstruction quality.. With limited regularization intensity controlled via adaptive gradients and decoupled from the photometric loss, a natural question arises as to whether existing regularization is appropriately effective. In practice, the regularization loss is thought to be controlled through hyperparameters, yet it remains unclear whether they provide sufficient flexibility. Moreover, can further decoupling enable more precise control over regularization itself? In this work, we focus on opacity and scaling regularization, which are closely tied to reconstruction quality in the 3DGS-MCMC framework (kheradmand20243d, ) and redundancy removal in the vanilla 3DGS framework (papantonakis2024reducing, ). Prior studies have shown that the densification stage often produces a large number of redundant primitives (liu2024maskgaussian, ), while the framework lacks a native mechanism to automatically remove them. Even with opacity regularization, existing pipelines usually rely on additional pruning operations or direct modifications to the densification process. However, if regularization can be well recomposed, is it possible to achieve redundancy removal without resorting to such extra pruning operations?
@@
Reanalyzing the complex coupling during 3DGS optimization, we takes a step to decouple it and recompose it into 3 effective components: Sparse Adam, Re-State Regularization and Decoupled Attribute Regularization.
@@
Based on the experiments and analysis, we propose the AdamW-GS with controllable attribute regularization and adopt it into vanilla 3DGS, 3DGS-MCMC and more variants. Experiments show that our optimization method yields better reconstruction quality and optimization efficiency, and significantly reducing redundancy in 3DGS without introducing additional pruning operations.
@@ 2 Related Work
“Coupling” in 3DGS Optimization: 3DGS commonly inherits the Adam optimizer from DNNs. However, its synchronous optimization causes primitives in invisible viewpoints to undergo overlooked update steps. Sparse Adam (mallick2024taming, ) introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. Regularization in 3DGS is also coupled with the photometric loss, akin to L2 weight regularization in adaptive gradient algorithms for DNNs. As analyzed in AdamW (loshchilov2017decoupled, ), the coupling of adaptive gradients in Adam makes the effective strength of regularization governed by the second moment, thereby motivating weight decay decoupling. A similar idea is reflected in (rota2025revising, ), which—though not explicitly designed for decoupling—replaces opacity reset in vanilla 3DGS with a constant opacity decay during densification, effectively constructing an AdamW-style optimizer. Nevertheless, this approach neglects the physical properties of opacity and the varying significance of individual primitives. The central contribution of this work is to further decouple and recompose 3DGS optimization and, through this process, reinterpret redundancy in 3DGS and reconstruction quality in 3DGS-MCMC from an optimization perspective.
Redundancy during Optimization: Typically, 3DGS requires careful treatment of the special process of primitive generation and pruning (or death (kheradmand20243d, ; zhu20253d, )). Current studies generally observe that optimization produces abundant redundant primitives, while pruning (in adaptive density control (kerbl20233d, )) alone remains insufficient to eliminate them. Four main strategies have been proposed to address this redundancy: (1) Densification refinement (fang2024mini, ; mallick2024taming, ; wang2025steepest, ); (2) Hand-crafted criterion (fan2024lightgaussian, ; niemeyer2025radsplat, ; hanson2025pup, ); (3) Learning-based pruning (lee2024compact, ; liu2024maskgaussian, ; zhang2024lp, ). More details can be found in Sec. F. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks (papantonakis2024reducing, ; lee2024compact, ; kheradmand20243d, ; liu2025deformable, ; svitov2024billboard, ).
@@ 3 Preliminary: 3DGS(-MCMC) and Adam
3DGS approximates the radiance field of a target scene using a set of Gaussian primitives parameterized by location, opacity (), scale (), and other attributes, and renders images from given viewpoints via alpha blending. Further details are provided in Appendix Sec. B.
Typically, the training pipeline can be divided into densification and pure optimization (P-Op). During densification, new primitives are generated to enhance scene representation while redundant ones are pruned (low-opacity primitives are removed) (kerbl20233d, ), or new primitives are sampled from existing primitives while dead primitives are reallocated to new locations (kheradmand20243d, ). After densification, there is a P-Op stage, during which only gradient propagation and attribute updates occur. In this training pipeline, photometric loss (left in Eq.1) is adopted with Adam optimizer. For 3DGS-MCMC, extra regularization loss in Eq. 1 is utilized to promote respawning.
@@
We omit the expression of bias correction procedure; in this paper, and denotes the primitive attribute in and gradient information respectively.
Consistent with the DNNs(goodfellow2016deep, ), Adam by default updates all attributes synchronously, including both the primitives visible under the current viewpoint and those that are not. In split of the 0 gradient, they are still included in the update step, during which their moment are rescaled and attributes are still updated. In this paper, updates of points in invisible viewpoints are referred to as Implicit Update, which are influenced by the previous moment.
@@ 4 Methodology
4.1 Characteristic of Sparse Adam
@@
To conduct a preliminary investigation, three experiments were performed on vanilla 3DGS: (GS1) training with the original Adam optimizer, (GS2) replacing Adam with Sparse Adam, and (GS3) employing Adam during the densification stage while applying Sparse Adam in the P-Op stage. Based on these comparisons in Table 1, we make the following observations: Observation 1 Sparse Adam is more stable: Since densification is governed by gradient magnitudes, the smaller number of primitives observed under Sparse Adam suggests that more primitives fail to meet the gradient threshold. Moreover, Sparse Adam holds on more active primitive, especially the GS3 in Table 1 (0.048 million dead primitives in Sparse Adam v.s. 0.232 million in Adam), which implies components in Adam are potential for pruning redundant primitives. Observation 2 Sparse Adam is less explorative: although it quickly drives primitives toward stability and results in fewer primitives overall, its performance degrades noticeably. Experiments on 3DGS-MCMC show the similar observation in some scenes that Sparse Adam makes smaller reallocated primitive number, and this phenomenon can be found in Figure 3 f.
@@ 4.2 Update-step decoupling: Re-State Regularization
Here, the subscript follows an asynchronous style different from synchronous style in Eq.2, meaning that each primitive has its own distinct . A detailed hyperparameters discussion are in Sec.H.2.
We also construct an implicit update to study its role. While implicit updates can be beneficial in some cases, they may introduce negative effects, potentially linked to redundancy (see Appendix Sec. G.4). In the experiment section of the mainbody, we retain implicit updates only as a tool to amplify attribute regularization under Sparse Adam to study the influence of regularization: when in Eq. 6 is large, maintaining implicit updates further strengthens regularization. Concretely, our Artificial Implicit Update (AIU) is defined as a uniform sampler that selects a random subset of invisible primitives for extra updates, while their moments remain unchanged in invisible viewpoints as in Sparse Adam.
@@ 4.3 Gradient Decoupling: Decoupled Attribute Regularization
We view attribute regularization as an essential component of optimization. In efficient 3DGS and related tasks, the widely-used opacity L1 regularization exemplifies this idea, analogous to weight regularization in DNNs, though with a different role. As discussed in Sec.3, our understanding of how regularization effectively constrain each primitive remains limited. Implicit updates may further exaggerate the effect of regularization: when in Eq.6 approaches or exceeds , the regularization term continues to influence the optimization via the adaptive gradient or implicit updates. The coupling of photometric and regularization losses in Eq.6 therefore raises a natural question: is the current regularization over- or under-effective?
First, we return to the factors actually govern the effect of the current attribute regularization loss (L1 loss with the gradient coupling). The role of the hyperparameter serves to maintain stability between losses, ensuring that the regularization remains relatively small at critical moments without interfering with the photometric loss. Relevant results are presented in Appendix Sec.G.1. Regarding the modulation by the moment, the regularization is activated when takes the majority or the moment is rescaled during synchronous optimization in Adam or reset in 3DGS-MCMC.
@@
Therefore, the first step is to decouple the gradients in the moment (Eq.6). This step is the same as AdamW, and readily yields Eq.7, which defines the moment update term containing only the gradient of the photometric loss . For , AdamW constructs updates that are not controlled by the second moment. This AdamW-style decoupling is effectively equivalent to a constant penalty on opacity (rota2025revising, ). The experiments are presented in Sec.G.2. Clearly, this approach is not optimal. Unlike DNNs, each attribute in 3DGS has physical meaning, and each primitive carries a distinct importance. Penalizing all primitives with the same strength fails to remove redundant primitives when the penalty is too weak, while overly strong penalties hinder optimization.
AdamW-style decoupling suggests that effective regularization assigns different penalties to different primitives. Such assignment, however, cannot be arbitrary: (i) it must not interfere with normal optimization, particularly in under-constructed regions ( is large); (ii) it must still fulfill its core function of suppressing overfitting when is small (srivastava2014dropout, ). This implies that regularization should adapt to the geometry of the data (pascanu2013revisiting, ), naturally linking back to adaptive gradient (duchi2011adaptive, ; kinga2015method, ) where adjustments are made via . The empirical success of vanilla regularization over the constant penalty further highlights its potential. Since we have already disentangled the moment in Eq. 7, now provides a more faithful estimate of the parameter space. Motivated by this, we introduce the regularization form ( is from Eq. 7). This design brings 4 benefits: (1) updates are preserved in under-constructed regions with large ; (2) when a primitive lies near a saddle point (small ), regularization facilitates escape, consistent with saddle-point analyses in (wang2025steepest, ) and long-axis primitive gradients in (zhang2024pixel, ); (3) regularization remains small in general but amplifies adaptively when it becomes influential; and (4) it enables explicit control over regularization via moment modulation, e.g., through RSR. Building on this rationale, we extend this approach and propose Eq. 8.
We maintain the loss balance through , and normalize the regularization scale by . Regularization aligned with improves stability while removing dependence on primitive number, which can vary by orders of magnitude across training, views, and scenes. keeps on the same order as maximum step size of Adam, matches the maximum observed in Figure 1 c–d, and extensive results confirm its robustness. Then, this recomposed regularization is referred to Decoupled Attribute Regularization (DAR). We provide detailed hyperparameter selection in Appendix Sec.H.2.
@@ 4.4 Recoupling: AdamW-GS
Building on Sec.4.1, 4.2, and 4.3, we propose AdamW-GS for better 3DGS optimization by recoupling Sparse Adam, RSR, and DAR. In detail, attributes of primitives are optimized asynchronously with Sparse Adam and supervised via photometric loss and DAR. At fixed training intervals, RSR uniformly samples primitives and rescales their moments via Eq. 5 to better activate regularization. The rescaled moment replaces the old one and directly participate in subsequent optimization.
We consider two DAR variants—opacity and scaling regularization—both of which are already employed in 3DGS-MCMC. Considering the bad transportation in vanilla 3DGS (jung2024relaxing, ), exploration in this work means better primitive movement. We directly use noise regularization in 3DGS-MCMC as an extra position regularization, which is similar to DAR, but controlled by opacity and primitive shape not the second moment. However considering the sensitivity to noise (jung2024relaxing, ), the noise is excluded for MipNerf360 indoors scenes with fewer primitives. To explore the effectiveness of exploration, we provide a contrast experiments with or without opacity reset(kerbl20233d, ). Opacity reset encourages exploration via lowering the opacity. Opacity Correction is adopted for cloning in densification as well, which is thought to be necessary(kheradmand20243d, ; rota2025revising, ). We adopt the similar operations for opacity like (rota2025revising, ).

### v3

@@ 1 Introduction
Novel view synthesis is a fundamental task in both computer graphics and computer vision. Owing to its highly parallelized design and efficient use of GPU, 3DGS (Kerbl et al., 2023) leverages explicit primitive-based representations to achieve significantly higher efficiency than implicit representation NeRF (Mildenhall et al., 2020), while still delivering competitive reconstruction quality. 3DGS sparks a wave of research and rapid extensions to various downstream applications (Matsuki et al., 2024; Chen et al., 2024; Dai et al., 2024).
@@
As a representation directly optimized via backpropagation, 3DGS inherits the Adam (Kinga et al., 2015) from DNNs, and in default attributes of all primitives are updated simultaneously within each iteration, which we call synchronous optimization in this work. Given the viewpoint relationship, an ideal scenario is that the optimization of primitives is guided by gradients from visible viewpoints in different directions. However, upon closer examination, we observe that even when primitives are invisible under a given viewpoint and their gradients are set to zero, the update-step coupling induced by Adam and synchronous optimization implicitly rescales the optimizer state and causes updating in these invisible viewpoints. From the perspective of optimization efficiency, (Mallick et al., 2024) introduces Sparse Adam, which restricts updates to primitives visible under viewpoints while filtering out those invisible with zero gradients, thereby enabling asynchronous optimization. Nevertheless, this optimizer has not been widely adopted, since its efficiency gains come at the expense of degraded performance. However, we think that such a viewpoint-related optimizer remains promising, with the current limitation lying in the insufficient recognition of update-step coupling.
@@
Furthermore, previous study (Loshchilov & Hutter, 2017) in DNNs points out the gradient coupling in Adam makes the regularization unstable. In current 3DGS, a variety of regularization losses (Yu et al., 2024) are introduced to improve reconstruction quality or mitigate redundancy111Redundancy is strictly defined as minimizing active primitives without sacrificing reconstruction quality.. With limited regularization intensity controlled via adaptive gradients and decoupled from the photometric loss, a natural question arises as to whether existing regularization is appropriately effective. In practice, the regularization loss is thought to be controlled through hyperparameters, yet it remains unclear whether they provide sufficient flexibility. Moreover, can further decoupling enable more precise control over regularization itself? In this work, we focus on opacity and scaling regularization, which are closely tied to reconstruction quality in the 3DGS-MCMC framework (Kheradmand et al., 2024) and redundancy removal in the vanilla 3DGS framework (Papantonakis et al., 2024). Prior studies have shown that the densification stage often produces a large number of redundant primitives (Liu et al., 2025b), while the framework lacks a native mechanism to automatically remove them. Even with opacity regularization, existing pipelines usually rely on additional pruning operations or direct modifications to the densification process. However, if regularization can be well recomposed, is it possible to achieve redundancy removal without resorting to such extra pruning operations?
@@
Reanalyzing the complex coupling during 3DGS optimization, we take a step to decouple it and recompose it into 3 effective components: Sparse Adam, Re-State Regularization and Decoupled Attribute Regularization.
@@
Based on the experiments and analysis, we propose the AdamW-GS with controllable attribute regularization and adopt it into vanilla 3DGS, 3DGS-MCMC and more variants. Experiments show that our optimization method yields better reconstruction quality and optimization efficiency, and significantly reduces redundancy in 3DGS without introducing additional pruning operations.
@@ 2 Related Work
“Coupling” in 3DGS Optimization: 3DGS commonly inherits the Adam optimizer from DNNs. However, its synchronous optimization causes primitives in invisible viewpoints to undergo overlooked update steps. Sparse Adam (Mallick et al., 2024) introduces asynchronous updates that improve efficiency, yet the insufficient recognition of update-step coupling results in degraded performance. Regularization in 3DGS is also coupled with the photometric loss, akin to L2 weight regularization in adaptive gradient algorithms for DNNs. As analyzed in AdamW (Loshchilov & Hutter, 2017), the coupling of adaptive gradients in Adam makes the effective strength of regularization governed by the second moment, thereby motivating weight decay decoupling. A similar idea is reflected in (Rota Bulò et al., 2025), which—though not explicitly designed for decoupling—replaces opacity reset in vanilla 3DGS with a constant opacity decay during densification, effectively constructing an AdamW-style optimizer. Nevertheless, this approach neglects the physical properties of opacity and the varying significance of individual primitives. The central contribution of this work is to further decouple and recompose 3DGS optimization and, through this process, reinterpret redundancy in 3DGS and reconstruction quality in 3DGS-MCMC from an optimization perspective.
Redundancy during Optimization: Typically, 3DGS requires careful treatment of the special process of primitive generation and pruning (or death (Kheradmand et al., 2024; Zhu et al., 2025)). Current studies generally observe that optimization produces abundant redundant primitives, while pruning (in adaptive density control (Kerbl et al., 2023)) alone remains insufficient to eliminate them. Four main strategies have been proposed to address this redundancy: (1) Densification refinement (Fang & Wang, 2024; Mallick et al., 2024; Wang et al., 2025); (2) Hand-crafted criterion (Fan et al., 2024; Niemeyer et al., 2025; Hanson et al., 2025); (3) Learning-based pruning (Lee et al., 2024; Liu et al., 2025b; Zhang et al., 2024a). More details can be found in Sec. H. (4) Optimization-related operations: Optimization together with additional opacity L1 regularization has become a common technique for redundancy removal, even extending beyond Efficient 3DGS tasks (Papantonakis et al., 2024; Lee et al., 2024; Kheradmand et al., 2024; Liu et al., 2025a; Svitov et al., 2024).
@@ 3 Preliminary: 3DGS(-MCMC) and Adam
3DGS approximates the radiance field of a target scene using a set of Gaussian primitives parameterized by location, opacity (), scale (), and other attributes, and renders images from given viewpoints via alpha blending. Further details are provided in Appendix Sec. D.
Typically, the training pipeline can be divided into densification and pure optimization (P-Op). During densification, new primitives are generated to enhance scene representation while redundant ones are pruned (low-opacity primitives are removed) (Kerbl et al., 2023), or new primitives are sampled from existing primitives while dead primitives are reallocated to new locations (Kheradmand et al., 2024). After densification, there is a P-Op stage, during which only gradient propagation and attribute updates occur. In this training pipeline, photometric loss (left in Eq.1) is adopted with Adam optimizer. For 3DGS-MCMC, extra regularization loss in Eq. 1 is utilized to promote respawning.
@@
We omit the expression of bias correction procedure; in this paper, and denote the primitive attribute in and gradient information respectively.
Consistent with the DNNs(Goodfellow et al., 2016), Adam by default updates all attributes synchronously, including both the primitives visible under the current viewpoint and those that are not. In spite of the 0 gradient, they are still included in the update step, during which their moment are rescaled and attributes are still updated. In this paper, updates of points in invisible viewpoints are referred to as Implicit Update, which are influenced by the previous moment.
@@ 4 Methodology
4.1 Characteristics of Sparse Adam
@@
To conduct a preliminary investigation, three experiments were performed on vanilla 3DGS: (GS1) training with the original Adam optimizer, (GS2) replacing Adam with Sparse Adam, and (GS3) employing Adam during the densification stage while applying Sparse Adam in the P-Op stage. Based on these comparisons in Table 1, we make the following observations: Observation 1 Sparse Adam is more stable: Since densification is governed by gradient magnitudes, the smaller number of primitives observed under Sparse Adam suggests that more primitives fail to meet the gradient threshold. Moreover, Sparse Adam holds on more active primitives, especially the GS3 in Table 1 (0.048 million dead primitives in Sparse Adam v.s. 0.232 million in Adam), which implies components in Adam are potential for pruning redundant primitives. Observation 2 Sparse Adam is less explorative: although it quickly drives primitives toward stability and results in fewer primitives overall, its performance degrades noticeably. Experiments on 3DGS-MCMC show the similar observation in some scenes that Sparse Adam makes smaller reallocated primitive number, and this phenomenon can be found in Figure 3 f.
@@ 4.2 Update-step decoupling: Re-State Regularization
Here, the subscript follows an asynchronous mode different from synchronous mode in Eq.2, meaning that each primitive has its own distinct . A detailed hyperparameters discussion is in Sec. J.2.
We also construct an implicit update to study its role. While implicit updates can be beneficial in some cases, they may introduce negative effects, potentially linked to redundancy (see Appendix Sec. I.4). In the experiment section of the mainbody, we retain implicit updates only as a tool to amplify attribute regularization under Sparse Adam to study the influence of regularization: when in Eq. 6 is large, maintaining implicit updates further strengthens regularization. Concretely, our Artificial Implicit Update (AIU) is defined as a uniform sampler that selects a random subset of invisible primitives for extra updates, while their moments remain unchanged in invisible viewpoints as in Sparse Adam.
@@ 4.3 Gradient Decoupling: Decoupled Attribute Regularization
We view attribute regularization as an essential component of optimization. In efficient 3DGS and related tasks, the widely-used opacity L1 regularization exemplifies this idea, analogous to weight regularization in DNNs, though with a different role. As discussed in Sec. 3, our understanding of how regularization effectively constrains each primitive remains limited. Implicit updates may further exaggerate the effect of regularization: when in Eq. 6 approaches or exceeds , the regularization term continues to influence the optimization via the adaptive gradient or implicit updates. The coupling of photometric and regularization losses in Eq. 6 therefore raises a natural question: is the current regularization over- or under-effective?
First, we return to the factors actually govern the effect of the current attribute regularization loss (L1 loss with the gradient coupling). The role of the hyperparameters serves to maintain stability between losses, ensuring that the regularization remains relatively small at critical moments without interfering with the photometric loss. Relevant results are presented in Appendix Sec. I.1. Regarding the modulation by the moment, the regularization is activated when takes the majority or the moment is rescaled during synchronous optimization in Adam or reset in 3DGS-MCMC.
@@
Therefore, the first step is to decouple the gradients in the moment (Eq. 6). This step is the same as AdamW, and readily yields Eq. 7, which defines the moment update term containing only the gradient of the photometric loss . For , AdamW constructs updates that are not controlled by the second moment. This AdamW-style decoupling is effectively equivalent to a constant penalty on opacity (Rota Bulò et al., 2025). The experiments are presented in Sec. I.2. Clearly, this approach is not optimal. Unlike DNNs, each attribute in 3DGS has physical meaning, and each primitive carries a distinct importance. Penalizing all primitives with the same strength fails to remove redundant primitives when the penalty is too weak, while overly strong penalties hinder optimization.
AdamW-style decoupling suggests that effective regularization assigns different penalties to different primitives. Such assignment, however, cannot be arbitrary: (i) it must not interfere with normal optimization, particularly in under-constructed regions ( is large); (ii) it must still fulfill its core function of suppressing overfitting when is small (Srivastava et al., 2014). This implies that regularization should adapt to the geometry of the data (Pascanu & Bengio, 2013), naturally linking back to adaptive gradient (Duchi et al., 2011; Kinga et al., 2015) where adjustments are made via . The empirical success of vanilla regularization over the constant penalty further highlights its potential. Since we have already disentangled the moment in Eq. 7, now provides a more faithful estimate of the parameter space. Motivated by this, we introduce the regularization form ( is from Eq. 7). This design brings 4 benefits: (1) updates are preserved in under-constructed regions with large ; (2) when a primitive lies near a saddle point (small ), regularization facilitates escape, consistent with saddle-point analyses in (Wang et al., 2025) and long-axis primitive gradients in (Zhang et al., 2024b); (3) regularization remains small in general but amplifies adaptively when it becomes influential; and (4) it enables explicit control over regularization via moment modulation, e.g., through RSR. Building on this rationale, we extend this approach and propose Eq. 8.
We maintain the loss balance through , and normalize the regularization scale by . Regularization aligned with improves stability while removing dependence on primitive number, which can vary by orders of magnitude across training, views, and scenes. keeps on the same order as maximum step size of Adam, matches the maximum observed in Figure 1 c–d, and extensive results confirm its robustness. Then, this recomposed regularization is referred to as Decoupled Attribute Regularization (DAR). We provide detailed hyperparameter selection in Appendix Sec. J.2.
@@ 4.4 Recoupling: AdamW-GS
Building on Sec. 4.1, 4.2, and 4.3, we propose AdamW-GS for better 3DGS optimization by recoupling Sparse Adam, RSR, and DAR. In detail, attributes of primitives are optimized asynchronously with Sparse Adam and supervised via photometric loss and DAR. At fixed training intervals, RSR uniformly samples primitives and rescales their moments via Eq. 5 to better activate regularization. The rescaled moment replaces the old one and directly participates in subsequent optimization.
We consider two DAR variants—opacity and scaling regularization—both of which are already employed in 3DGS-MCMC. Considering the bad transportation in vanilla 3DGS (Jung et al., 2024), exploration in this work means better primitive movement. We directly use noise regularization in 3DGS-MCMC as an extra position regularization, which is similar to DAR, but controlled by opacity and primitive shape not the second moment. However considering the sensitivity to noise (Jung et al., 2024), the noise is excluded for MipNerf360 indoors scenes with fewer primitives. To explore the effectiveness of exploration, we provide contrast experiments with or without opacity reset(Kerbl et al., 2023). Opacity reset encourages exploration via lowering the opacity. Opacity Correction is adopted for cloning in densification as well, which is thought to be necessary (Kheradmand et al., 2024). We adopt the similar operations for opacity as (Rota Bulò et al., 2025).